OpenStudy (anonymous):

kevin and his dad were hiking along the florida national scenic trail. their goal for the day was to hike 8 1/2 miles after 2 hours of hiking they had hiked 4 3/5miles. how many miles do they have left to hike in order to complete their goal of 8 1/2 miles? a. 3 1/2 b. 3 9/10 c. 4 2/3 d. 4 2/4

hero (hero):

Hint: After two hours, they hiked 4 3/5 miles. They still have x miles to go to reach 8 1/2 miles. Therefore, the equivalent expression is: 4 3/5 miles + x miles = 8 1/2 miles or simply 4 3/5 + x = 8 1/2 Solve for x

OpenStudy (anonymous):

4 3/5 + x =8 1/2

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

can you recheck the problem? is it 81/2 miles and then 43/5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or is it answer b that your wrote wrong was it 319/10?

OpenStudy (anonymous):

3 1/2

OpenStudy (anonymous):

are the answers in miles as the unit?

hero (hero):

2nd Hint x = 8 1/2 - 4 3/5

OpenStudy (anonymous):

i did that

hero (hero):

What did you get?

OpenStudy (anonymous):

3 5/10 but i need to reduce it

hero (hero):

Actually, here's a little trick: 1/2 = .5 3/5 = .6 so... x = 8.5 - 4.6 = 3.9 .9 = 9/10 so x = 3 9/10

OpenStudy (anonymous):

it b

OpenStudy (anonymous):

can you explain more please

hero (hero):

Yes, correct. It is b. I didn't even see that above list of answer choices.

hero (hero):

What exactly is it that you don't understand?

OpenStudy (anonymous):

Hm.....like so if u get a prob that says how many many miles do they have left to get the goal. I do it sometimes by adding i always learned it that way but in this prob i didnt get the answer that so, do u always have to subtract?

hero (hero):

Let's start with something simple. Suppose you want to have 8 apples, but you only have three so far. How many apples do you need to get 8?

hero (hero):

What grade are you in @partyrainbow276 ?

OpenStudy (anonymous):

So, i would use the choices and add them like..... 3+5 A. 5 B. 6 C. 14 So... i would put A. D. 20

OpenStudy (anonymous):

y?

hero (hero):

Because I need to know how to explain things on a level where you can understand them.

OpenStudy (anonymous):

6th grade?

hero (hero):

I see

hero (hero):

But the thing is, we didn't know it was 5 when we first started. It was some unknown number that we had to add to 3 to get 8. IIn Algebra, an unknown number is represented by a variable, usually x. In this case, it would be 3 + x = 8 Now you would isolate x on one side of the equation by solving for x: x = 8 - 3 x = 5 So yes, x = 5 You must understand this process in order to solve the problems you are working with. That being said, usually different problems will have different algebraic statements associated with them. Sometimes the expression will involve addition, sometimes subtraction, sometimes, multiplication or division.

hero (hero):

The way that the apple problem was setup is similar to the hiking problem. With objects, You have a certain number of something, but you need the rest of it to get a total amount. With distance, you went so far, but you need to get the rest of the way to complete the trip.

OpenStudy (anonymous):

kay kay ^-^ Thank uuuuuuuuuuu!! so....... MUCH